OpenStudy (anonymous):

Simplify √-121 a. -11i b. 11i c. -11 d. 11

OpenStudy (anonymous):

R=Radical -1(R)=?

OpenStudy (anonymous):

11i

OpenStudy (anonymous):

This problem is from Complex Numbers ALG 2 - T1 Rational, Complex, and Polynomials

OpenStudy (anonymous):

\[\sqrt{-121} = \sqrt{-1}\times \sqrt{121} = i \sqrt{11}\]